Erie Company manufactures a mobile fitness device called the Jogging Mate. The company uses standards to control its costs. The labor standards that have been set for one Jogging Mate are as follows:

Standard
Hours
Standard Rate
per Hour
Standard
Cost
27 minutes $6.20 $2.79

During August, 9,060 hours of direct labor time were needed to make 19,000 units of the Jogging Mate. The direct labor cost totaled $54,360 for the month.

Required:

1. What is the standard labor-hours allowed (SH) to makes 19,000 Jogging Mates?

2. What is the standard labor cost allowed (SH × SR) to make 19,000 Jogging Mates?

3. What is the labor spending variance?

4. What is the labor rate variance and the labor efficiency variance?

5. The budgeted variable manufacturing overhead rate is $4.30 per direct labor-hour. During August, the company incurred $43,488 in variable manufacturing overhead cost. Compute the variable overhead rate and efficiency variances for the month.

(For requirements 3 through 5, indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do not round intermediate calculations.)
Answer #1

1.Standard Hours Allowed = Number of Units*Hours per unit

= 19,000*27/60 = 8,550 Hours

2.Standard Labor cost = 8,550*6.20 = $53,010

3.Labor Spending Variance = Standard Cost – Actual Cost

= 53,010 – 54,360

= 1,350 U

4.Labor Rate Variance = (Standard rate – Actual Rate)*Actual Hours

= (6.20 – Actual rate)*9,060

= 6.20*9060 – 54,360

= $1,812 F

Efficiency = (Standard Hours – Actual Hours)*Standard Rate

= (8,550-9,060)*6.20

= $3,162 U

5.Variable overhead rate variance = (Standard rate – Actual Rate)*Actual Hours

= (4.30 – Actual rate)*9060

= 4.30*9060 – 43,488

= $4,530 U

Efficiency = (Standard Hours – Actual Hours)*Standard Rate

= (8550-9060)*4.30

= $2,193 U